This year, Mesa Inc.’s before-tax income was $11,377,000. It paid $539,000 income tax to Minnesota and $457,000 income tax to Illinois.

a. Compute Mesa’s federal income tax.

b. What is Mesa’s tax rate on its income?
Answer #1

Total state income tax = M’s tax + I’s tax

                                    = 539,000 + 457,000

                                    = 996,000

This is to be deducted from before-tax income for getting “taxable income for Federal tax purpose”.

Taxable income for Federal tax purpose = Before-tax income – Total state income tax

                                                               = 11,377,000 – 996,000

                                                                 = 10,381,000

This year corporate tax rate is 21%.

a.

Federal income tax = Taxable income for Federal tax purpose × Corporate tax rate

                                    = 10,381,000 × 21%

                                    = 2,180,010 (Answer)

b.

Total tax = 996,000 + 2,180,010 = 3,176,010

Tax rate = (Total tax / Before-tax income) × 100

            = (3,176,010 / 11,377,000) × 100

            = 27.92% (Answer)